OpenStudy (anonymous):

What is the x-coordinate of the x-intercept of the line that passes through points (-5, 3) and (1, -3)?

OpenStudy (mokeira):

First you need to form the equation of a line with those two points. Do you know how to form equation of a line?

OpenStudy (mokeira):

@mo12us

OpenStudy (anonymous):

no i'm lost can you walk me through it

OpenStudy (anonymous):

@Mokeira

OpenStudy (mokeira):

First step, find the slope \[m=\frac{ \Delta y }{ \Delta x }\] \[m=\frac{ 3--3 }{ -5-1 }\] \[m=\frac{ 6 }{ -6 }\] \[m=-1 \] Now introduce an unknown point (x,y) Then form equation of that line \[-1=\frac{ y-3 }{ x--5 }\] \[-1=\frac{ y-3 }{ x+5 }\] Cross multiply \[-1(x+5)=y-3\] \[-x-5=y-3\] \[y=-x-2\] The x coordinate is therefore -1 and to find x intercept,y=0 now substitute

OpenStudy (mokeira):

hope you understood

OpenStudy (anonymous):

thanks, sorry for the delay

OpenStudy (anonymous):

@Mokeira

OpenStudy (mokeira):

its cool